OpenStudy (ethan01):

Select the correct difference. -3z 7 - (-5z 7) A.2z7 B.8z14 C.-2z7 D.2z

OpenStudy (ethan01):

thanks

OpenStudy (ethan01):

Select the correct difference. 5d 2 + 4d - 3 less 2d 2 - 3d + 4 3d2 + 7d - 7 3d2 - 7d - 7 7d2 + 7d - 7 7d2 + d + 1

OpenStudy (supersmart1001):

is that it?

OpenStudy (ethan01):

yup

OpenStudy (supersmart1001):

ok

TheSmartOne (thesmartone):

Add like terms. Hint: \(\Large -3z^7 - (-5z^7) = -3z^7 + 5z^7 = (-3 + 5)z^7 = ??\)